PLease help!!!!! I cannot get anyone to answer this correctly.... Please write legibly and provide answers for all parts of this one question:

Consider the following contingency table.

B Bc
A 19 25
Ac 31 25


a. Convert the contingency table into a joint probability table. (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.)

B Bc Total
A
Ac
Total


b. What is the probability that A occurs? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.)
  

c. What is the probability that A and B occur? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.)
  

d. Given that B has occurred, what is the probability that A occurs? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.)
  

e. Given that Ac has occurred, what is the probability that B occurs? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.)
Answer #1
B Bc Total
A 0.19 0.25 0.44
Ac 0.31 0.25 0.56
Total 0.50 0.50 1

b) P(A)= 0.4400

c)P(A λ B) = 0.1900

d) P(A|B) =0.19/0.50= 0.3800

e) e)P(B|Ac) = P(Ac)

=0.31/0.56= 0.5536